3. Using Method of Sections determine the forces in members BC, CF, and GF. 200 TL 200 lb 200 lb Scale: 1 unit = 5 ft
Answer #1

A Ś Bisc sost E Faso Anao { fj=0 zools zoolb zoo16 Ay+ Ey a boolb o EME=0 + Ayx 20 - 200% 15-200X10-200x5 =0 Ay = 300lb, Ey& Me zo + 300X10 – 2005 – fer x = 0 . [fra = 400lb Tension from ③ - FBC = -400 +141.42 cos 45 FBC = -300 16 T compression.